An aspirin tablet contains 325 mg of the active ingredient, acetylsalicylic acid. Acetylsalicylic acid has the chemical formula C9H8O4. Calculate the number of moles of acetylsalicylic acid in this tablet of aspirin.
1
Expert's answer
2020-03-04T08:00:10-0500

Mr(C9H8O4) = 12*9+1*8+16*4=180.

325/1000*180=0.00180555556 moles